Please help!
Barn painting. Jasmine can paint a certain barn by herself in x days. Her helper Mindy can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x=5

Answer by stanbon(48558)

Barn painting. Jasmine can paint a certain barn by herself in x days. Her helper Mindy can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x=5
---------------------
Jasmine DATA:
time = x days/job ; rate = 1/x job/day
------------------------------------------------
Minday DATA:
time = 2x days/job ; rate = 1/2x job/day
-------------------------------------------------
Write a rational expression for the fraction of the barn that they complete in one day by working together.
---
One day : (1/x) + 1/(2x) = (2x+x)/2x^2 = 3x/2x^2 = 3/2x of the job
Evaluate the expression for x=5
If x = 5 : 3/(2x) = 3/10 of the job
=======================================
Cheers,
Stan H.